### v2

# Quantitative estimates for singularity for conjugate equations driven by linear fractional transformations

Abstract.
We consider the conjugate equation driven by two families of finite maps on the unit interval satisfying a compatibility condition. This framework contains de Rham’s functional equations. We give sufficient conditions for singularity of the solution with quantitative estimates in the case that the equation is driven by non-affine maps, in particular, linear fractional transformations.
1. Introduction
Let $N\geq 2$ and let $\mathcal{I}_{N}:=\{0,1,\cdots,N-1\}$. Let $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ be two families of $N$ maps on $[0,1]$ satisfying a compatibility condition. We consider the following conjugate equation for $\varphi:[0,1]\to[0,1]$:
$g_{i}\circ\varphi=\varphi\circ f_{i},\ i\in\mathcal{I}_{N}.$ (1.1)
See Section 2 below for the definition of the compatibility condition. An alternative expression is given by
$\varphi(x)=\begin{cases}g_{0}(\varphi(f_{0}^{-1}(x)))&0\leq x\leq f_{0}(1)\\ \vdots&\\ g_{N-1}(\varphi(f_{N-1}^{-1}(x)))&f_{N-1}(0)\leq x\leq 1\end{cases}.$ (1.2)
This is known as de Rham’s functional equation [7, 8]. This functional equation is generally fractal in nature, and solutions of (1.1) can be singular functions.
Assume that $N=2$, $10<p<$ 1, $f_{0}(x)=x/2$, $f_{1}(x)=(x+1)/2$, $g_{0}(x)=px$ and $g_{1}(x)=(1-p)x+p$. Then, (1.2) is expressed as
$\varphi(x)=\begin{cases}p\varphi(2x)&0\leq x\leq 1/2\\ (1-p)\varphi(2x-1)+p&1/2\leq x\leq 1\end{cases}.$
There exists a unique continuous strictly increasing solution $\varphi_{p}$ for this equation, and $\varphi_{p}$ is singular if $p\neq 1/2$ and $\varphi_{p}(x)=x$ if $p=1/2$. $\varphi_{p}$ is called the Lebesgue singular function. This function is also a solution of the functional equation by Matkowski and Wesołowski [16, 24, 25].
Minkowski’s question-mark function [23] is also obtained by a solution of (1.1) for the case that $(f_{0}(x),f_{1}(x),g_{0}(x),g_{1}(x))=\left(x/(x+1),1/(2-x),x/2,(x+1)/2\right)$. It can also be expressed in the form
$\varphi(x)=\begin{cases}\frac{1}{2}\varphi(\frac{x}{1-x})&0\leq x\leq 1/2\\ \frac{1}{2}\varphi\left(2-\frac{1}{x}\right)+\frac{1}{2}&1/2\leq x\leq 1\end{cases}.$
This function has appeared in various areas of mathematics such as number theory and dynamical systems. The real analytic properties of this function have been considered ([15, 19, 21]). The conjugate equation of the de Rham type has been investigated by many authors ([2, 4, 6, 11, 12, 13, 17, 28, 29, 30, 32]). We refer the reader to [27, Subsection 3.4] for a more detailed review of de Rham’s functional equations.
The main purpose of this paper is to investigate some properties of the solution of (1.1). Specifically, we give sufficient conditions for the solution to be singular. Since $\varphi$ is an increasing function on $[0,1]$, we deal with the probability measure $\mu_{\varphi}$, which is the Borel probability measure on $[0,1]$ whose distribution function is $\varphi$. We consider the case that both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are non-affine. This case is more difficult to analyze than the case that both $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are affine.
It is interesting to know a necessary and sufficient condition of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ for the singularity. This topic has been considered in [2, 3, 10, 13, 18, 25]. In [27], the author dealt with the case that $\{g_{i}\}_{i}$ are linear fractional transformations and $\{f_{i}\}_{i}$ are affine. Our results, Theorems 3.2 and 3.3 below, deal with the case that $\{f_{i}\}_{i}$ are not necessarily affine. We do not assume the differentiability of $\{f_{i}\}_{i}$. For the proof, we use a martingale approach as in [27], but here we need more delicate quantitative estimates.
The structure of this paper is as follows. In Section 2, we give some terminology and basic results of the conjugate equation on the unit interval. In Section 3, we deal with the singularity. Finally, we give an example in Section 4.
2. Preliminaries
2.1. Existence and uniqueness
We first discuss the existence and uniqueness of the solution of the conjugate equation (1.1).
We say that a family of maps $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ is a compatible system or satisfies a compatibility condition if they are strictly increasing continuous functions such that $h_{0}(0)=0$, $h_{i-1}(1)=h_{i}(0)$, $11\leq i\leq N-$ 1, $h_{N-1}(1)=1$, and, a compatible system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system on $[0,1]$ if
$\mathcal{D}_{h}:=\{h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j):i_{1},\cdots i_{n}\in\mathcal{I}_{N},j\in\{0,1\}\}$
is dense in $[0,1]$.
Lemma 2.1.
Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be a D-system on $[0,1]$. Then, for every $x\in[0,1]$, there exists an infinite sequence $(i_{n})_{n}$ of $\mathcal{I}_{N}$ such that
$x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1).$ (2.1)
If $x\in\mathcal{D}_{h}\setminus\{0,1\}$, then, there exist two such infinite sequences. Otherwise, there exists exactly one such infinite sequence.
Proof.
We first show that for every infinite sequence $(i_{n})_{n}$,
$\lim_{n\to\infty}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)|=0.$ (2.2)
Assume this fails. By the definition of $\{h_{i}\}_{i}$, $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right)_{n}$ is increasing and $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right)_{n}$ is decreasing. Since $\mathcal{D}_{h}$ is dense in $[0,1]$, there exists $k_{1},\cdots,k_{N}$ and $j\in\{0,1\}$ such that
$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)<h_{k_{1}}\circ\cdots\circ h_{k_{N}}(j)$
$<\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
This cannot occur. Hence (2.2) holds.
Case 1. Assume $x\in\mathcal{D}_{h}\setminus\{0,1\}$. Assume that $x=h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)$. Since $h_{0}(0)=0$ and $x>0$, we can assume that $i_{n}\geq 1$. Let $i_{k}:=0$ for every $k\geq n+1$. Then, (2.1) holds. By replacing $i_{n}$ with $i_{n}-1$ and let $i_{k}:=N-1$ for every $k\geq n+1$, then, (2.1) holds. Assume that $x=h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)$. Since $h_{N-1}(1)=1$ and $x<1$, we can assume that $i_{n}\leq N-2$. Let $i_{k}:=N-1$ for every $k\geq n+1$. Then, (2.1) holds. By replacing $i_{n}$ with $i_{n}+1$ and let $i_{k}:=0$ for every $k\geq n+1$, then, (2.1) holds.
Case 2. Assume $x\in\{0,1\}$. Assume $x=0$. Since each $h_{i}$ is strictly increasing, there is no other choice than $i_{k}=0$ for every $k\geq 1$ for (2.1). Assume $x=1$. Then, there is no other choice than $i_{k}=N-1$ for every $k\geq 1$ for (2.1).
Case 3. Assume $x\notin\mathcal{D}_{h}$. Then, for each $n\geq 1$, there exists a unique finite sequence $(i_{n,1},\cdots,i_{n,n})$ such that
$h_{i_{n,1}}\circ\cdots\circ h_{i_{n,n}}(0)<x<h_{i_{n,1}}\circ\cdots\circ h_{i_{n,n}}(1).$
Then, $i_{n-1,k}=i_{n,k}$ for every $k\leq n-1$. Let $i_{k}:=i_{n,k}$. By (2.2), (2.1) holds. If (2.1) holds, then, for every $N\geq 1$,
$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
Hence,
$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)<x<h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
Hence the uniqueness holds. ∎
Remark 2.2.
The existence and uniqueness of the bounded solutions of conjugate equations on more general spaces than $[0,1]$ have been discussed in [6, 26, 28, 30].
Lemma 2.3.
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are D-systems on $[0,1]$. Then, there exists a unique continuous strictly increasing solution of (1.1).
Hereafter, we let $\mu_{\varphi}$ be the continuous probability measure on $[0,1]$ whose distribution function is $\varphi$. This is the central object considered in this paper.
Proof.
We first see that
$\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(j)\right)=g_{i_{1}}\circ\cdots\circ g_{i_{n}}(j),\ i_{1},\cdots,i_{n}\in\mathcal{I}_{N},\ j\in\{0,1\},$ (2.3)
by induction in $n$. Hence, $\varphi(\mathcal{D}_{f})=\mathcal{D}_{g}$.
Step 1. We define $\varphi(x)$ for $x\notin\mathcal{D}_{f}$. By the assumption, there exists an infinite sequence $(i_{n})_{n}$ of $\mathcal{I}_{N}$ such that $x\in(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1))$. Since $\left[g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0),g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)\right]$ is decreasing with respect to $n$,
$\bigcap_{n\geq 1}\left[g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0),g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)\right]=\left[\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0),\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)\right].$
By the density of $\mathcal{D}_{g}$ in $[0,1]$, we can apply (2.2), and obtain that
$\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1).$ (2.4)
We let $\varphi(x):=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)$.
Step 2. We show $\varphi$ is strictly increasing on $[0,1]$. By (2.3), $\varphi$ is strictly increasing on $\mathcal{D}_{f}$. Let $x<y$. By (2.4), for some $n$,
$f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0)\leq x\leq f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)<f_{j_{1}}\circ\cdots\circ f_{j_{n}}(0)\leq y\leq f_{j_{1}}\circ\cdots\circ f_{j_{n}}(1).$
By this and Step 1,
$\varphi(x)\leq g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)<g_{j_{1}}\circ\cdots\circ g_{j_{n}}(0)\leq\varphi(y).$
Step 3. We show that $\varphi$ is continuous on $[0,1]$. Assume that there is a discontinuous point $x_{0}\in(0,1)$ of $\varphi$. Then,
$b_{0}:=\lim_{x\to x_{0}+0}\varphi(x)>\lim_{x\to x_{0}-0}\varphi(x)=:a_{0}.$
By Step 2, $\varphi([0,1])\cap(a_{0},b_{0})=\emptyset$, in particular, $\mathcal{D}_{g}\cap(a_{0},b_{0})=\emptyset$. This contradicts the density of $\mathcal{D}_{g}$ in $[0,1]$. The same argument is applied to the case that $x_{0}\in\{0,1\}$. ∎
It is natural to consider the inverse function of the solution $\varphi$. We easily see that
Lemma 2.4.
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are D-systems on $[0,1]$. Let $\varphi$ be a unique solution of (1.1). Then, the inverse function $\varphi^{-1}$ of $\varphi$ is well-defined and satisfies the following conjugate equation
$\varphi^{-1}\circ g_{i}=f_{i}\circ\varphi^{-1},\ i\in\mathcal{I}_{N},$
which is equivalent with (1.1) where $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are exchanged.
By [4, Proposition 3.1], if $\varphi$ is singular, then, $\varphi^{-1}$ is also singular.
2.2. Weak contraction
Hereafter, we assume more additional conditions for the D-systems $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$. We say that a function $h:[0,1]\to[0,1]$ is a contraction on $[0,1]$ if
$\textup{Lip}(h):=\sup_{x\neq y}\left|\frac{h(x)-h(y)}{x-y}\right|<1.$
$\textup{Lip}(h)$ is called the Lipschitz constant of $h$. We say that a function $h:[0,1]\to[0,1]$ is a weak contraction on $[0,1]$ if
$\phi_{h}(t):=\lim_{s\to t+0}\sup\left\{|h(x)-h(y)|:|x-y|\leq s\right\}<t,\ \ t>0.$
The following is well-known. See [5, 14].
Lemma 2.5.
Let $h$ be a weak contraction. Then, $\lim_{n\to\infty}\phi_{h}^{n}(t)=0$ for every $t>0$ and $|h(x)-h(y)|\leq\phi_{h}(|x-y|)$ for $x,y\in[0,1]$.
Proof.
Let
$\omega_{h}(s):=\sup\left\{|h(x)-h(y)|:|x-y|\leq s\right\}.$
Since this function is increasing, $\phi_{h}$ is right-continuous and increasing on $[0,1]$. Hence, $|h(x)-h(y)|\leq\lim_{s\to|x-y|+0}\phi_{h}(s)=\phi_{h}(|x-y|)$.
By the assumption, for every $t>0$, $(\phi_{h}^{n}(t))_{n}$ is strictly decreasing. There exists the limit $\phi^{\infty}_{h}(t):=\lim_{n\to\infty}\phi_{h}^{n}(t)$. By the right-continuity of $\phi_{h}$, $\phi_{h}\left(\phi^{\infty}_{h}(t)\right)=\phi^{\infty}_{h}(t)$. Hence, $\phi^{\infty}_{h}(t)=0$. ∎
Lemma 2.6.
Let $\{h_{i}\}_{i}$ be a compatible system on $[0,1]$ such that each $h_{i}$ is a weak contraction on $[0,1]$. Then, $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system.
We say that a compatible system $\{h_{i}\}_{i}$ on $[0,1]$ is an (weakly) contractive system on $[0,1]$ if each $h_{i}$ is an (weak) contraction on $[0,1]$. Every weakly contractive system is a D-system.
Proof.
Let $\phi:=\max_{i\in\mathcal{I}_{N}}\phi_{h_{i}}$. Then, $\phi$ is right-continuous [^1] The assumption that $N$ is finite would be important. on $[0,1]$, $\phi(t)<t$ for every $t>0$, and $\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\phi(|x-y|)$, $x,y\in[0,1]$. Then, we see that $\lim_{n\to\infty}\phi^{n}(t)=0$ for every $t>0$. Hence,
$\left|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right|\leq\phi^{n}(1).$
$\lim_{n\to\infty}\max_{i_{1},\cdots,i_{n}\in\mathcal{I}_{N}}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)|=0.$
By the definition, for each $n\geq 1$,
$[0,1]=\bigcup_{i_{1},\cdots,i_{n}\in\mathcal{I}_{N}}\left[h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0),h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right],$
and their interiors are disjoint. ∎
2.3. Linear fractional transforms
Definition 2.7.
We say that a family of functions $\{h_{i}\}_{i}$ is an LF system on $[0,1]$ if $h_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, for each $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers such that
(1) (well-defined) $c_{i}+d_{i}>0$, $d_{i}>0$, $i\in\mathcal{I}_{N}$.
(2) (strictly increasing, weak contraction)
$0<a_{i}d_{i}-b_{i}c_{i}\leq\min\{d_{i}^{2},(c_{i}+d_{i})^{2}\},\ i\in\mathcal{I}_{N}.$
(3) (compatibility) $b_{0}=0$, $\dfrac{a_{i}+b_{i}}{c_{i}+d_{i}}=\dfrac{b_{i+1}}{d_{i+1}}$, $i\in\mathcal{I}_{N}$, where we let $\dfrac{b_{N}}{d_{N}}:=1$.
For condition (2), we may consider a stronger condition that
(2’) $0<a_{i}d_{i}-b_{i}c_{i}<\min\{d_{i}^{2},(d_{i}+c_{i})^{2}\},\ i\in\mathcal{I}_{N}$.
Since $h_{i}$ is invariant if we multiply a positive constant by each of $a_{i},b_{i},c_{i}$ and $d_{i}$, we may assume that $d_{i}=1$, or, $a_{i}d_{i}-b_{i}c_{i}=1$.
Lemma 2.8.
Every LF system is a weakly contractive system. If (2’) holds instead of (2), then, it is a contractive system.
Proof.
Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF-system. We assume that $d_{i}=1$ for every $i$. By Lemma 2.6, it suffices to show that each $h_{i}$ is a weak contraction. We first remark that
$|h_{i}(x)-h_{i}(y)|=\frac{a_{i}-b_{i}c_{i}}{(c_{i}x+1)(c_{i}y+1)}|x-y|.$
We consider three cases: (i) $c_{i}>0$, (ii) $c_{i}<0$, and (iii) $c_{i}=0$.
(i) Assume $c_{i}>0$. Then, $a_{i}-b_{i}c_{i}\leq 1$ and
$|h_{i}(x)-h_{i}(y)|\leq\frac{a_{i}-b_{i}c_{i}}{c_{i}|x-y|+1}|x-y|\leq\frac{|x-y|}{c_{i}|x-y|+1}.$
Hence,
$\omega_{h_{i}}(s)\leq\frac{s}{c_{i}s+1},\ \ s\geq 0.$
By this and $c_{i}>0$,
$\phi_{h_{i}}(t)\leq\frac{t}{c_{i}t+1}<t,\ t>0.$
(ii) Assume $c_{i}<0$. Then, $a_{i}-b_{i}c_{i}\leq(c_{i}+1)^{2}$ and
$|h_{i}(x)-h_{i}(y)|\leq\frac{a_{i}-b_{i}c_{i}}{(c_{i}+1)(c_{i}(1-|x-y|)+1)}|x-y|\leq\frac{(1+c_{i})|x-y|}{-c_{i}|x-y|+1+c_{i}}.$
Hence,
$\omega_{h_{i}}(s)\leq\frac{(1+c_{i})s}{-c_{i}s+1+c_{i}},\ \ s\geq 0.$
By this and $c_{i}<0$,
$\phi_{h_{i}}(t)\leq\frac{(1+c_{i})t}{-c_{i}t+1+c_{i}}<t,\ t>0.$
(iii) Assume $c_{i}=0$. Then, $0<a_{i}\leq 1$ and $|h_{i}(x)-h_{i}(y)|=a_{i}|x-y|$. By the assumption for $a_{i},b_{i},c_{i}$, $h_{j}(1)>h_{j}(0)$ for every $j$. Hence, by $N\geq 2$, $a_{i}=h_{i}(1)-h_{i}(0)<1$. ∎
Remark 2.9.
The technique used above is related to a characterization of Rakotch contractions. See [20, Lemma 2.10 (iii)].
Let $(\mathcal{I}_{N})^{*}:=\cup_{n\in\mathbb{N}}(\mathcal{I}_{N})^{n}$. We denote the length of $\sigma\in(\mathcal{I}_{N})^{*}$ by $|\sigma|$. For an LF system $\{h_{i}\}_{i}$ and $\sigma=(\sigma_{1},\cdots,\sigma_{n})\in(\mathcal{I}_{N})^{*}$, we let
$h_{\sigma}(x)=h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
Lemma 2.10.
Let $\{h_{i}\}_{i}$ be an LF system. Then, there exists $c\in(0,1)$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
$h_{\sigma}(1)-h_{\sigma}(0)\geq c^{|\sigma|}.$
Proof.
Let $A_{i}:=\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}$, $i\in\mathcal{I}_{N}$. We assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i$. Let $(i_{n})_{n}\in(\mathcal{I}_{N})^{\mathbb{N}}$. Let
$\begin{pmatrix}p_{n}&q_{n}\\ r_{n}&s_{n}\end{pmatrix}=A_{i_{1}}\cdots A_{i_{n}},\ n\geq 1.$
Let $C_{1}:=\max\{|a_{i}|,|b_{i}|,|c_{i}|,|d_{i}|:i\in\mathcal{I}_{N}\}$. Then, for every $n$, $|r_{n+1}|+|s_{n+1}|\leq C_{1}(|r_{n}|+|s_{n}|)$. Hence, $|r_{n}|+|s_{n}|\leq 2C_{1}^{n}$ for every $n$.
Assume that $\sigma=(i_{1},\cdots,i_{n})$. Then, $|\sigma|=n$ and
$h_{\sigma}(1)-h_{\sigma}(0)=\frac{1}{s_{n}(r_{n}+s_{n})}\geq\frac{1}{(|r_{n}|+|s_{n}|)^{2}}\geq\frac{1}{4|C_{1}|^{2n}}.$
We give an upper bound for $h_{\sigma}(1)-h_{\sigma}(0)$.
Lemma 2.11.
Let $\{h_{i}\}_{i}$ be an LF system. Then, there exists $C>0$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
$h_{\sigma}(1)-h_{\sigma}(0)\leq\frac{C}{|\sigma|}.$
Proof.
We first remark that
$(c_{i}x+d_{i})(c_{i}y+d_{i})\geq 1+|c_{i}||x-y|,\ x,y\in[0,1].$ (2.5)
We assume that $x\leq y$. Assume $c_{i}>0$. Then,
$c_{i}y+d_{i}\geq c_{i}x+d_{i}\geq\min_{x\in[0,1]}c_{i}x+d_{i}=\min\{d_{i},c_{i}+d_{i}\}=d_{i}\geq 1.$
Hence,
$(c_{i}x+d_{i})(c_{i}y+d_{i})=(c_{i}x+d_{i})^{2}+c_{i}(y-x)(c_{i}x+d_{i})\geq 1+c_{i}(y-x).$
Assume $c_{i}<0$. Then,
$(c_{i}x+d_{i})(c_{i}y+d_{i})=(c_{i}y+d_{i})^{2}-c_{i}(y-x)(c_{i}x+d_{i})\geq 1-c_{i}(y-x).$
If $c_{i}=0$, then, $(c_{i}x+d_{i})(c_{i}y+d_{i})=d_{i}^{2}\geq 1=1+|c_{i}||x-y|$. Thus we see (2.5).
By (2.5),
$|h_{i}(x)-h_{i}(y)|\leq\frac{|x-y|}{1+|c_{i}||x-y|}.$
If $c_{i}=0$, then,
$|h_{i}(x)-h_{i}(y)|=\frac{a_{i}}{d_{i}}|x-y|.$
Since $h_{i}(1)-h_{i}(0)<1$, $a_{i}<d_{i}$. Let $C_{1}:=\min\{|c_{i}|:c_{i}\neq 0\}$ and $C_{2}:=\max\left\{\dfrac{a_{i}}{d_{i}}:c_{i}=0\right\}$. Then, $C_{1}>0$, $C_{2}<1$ and
$\max_{i}|h_{i}(x)-h_{i}(y)|\leq\max\left\{\frac{|x-y|}{1+C_{1}|x-y|},C_{2}|x-y|\right\}.$
We can take $C_{3}\in(0,C_{1})$ such that
$\max_{i}|h_{i}(x)-h_{i}(y)|\leq\frac{|x-y|}{1+C_{3}|x-y|},\ x,y\in[0,1].$
Let $a_{n}:=\max\left\{h_{\sigma}(1)-h_{\sigma}(0):\sigma\in(\mathcal{I}_{N})^{*},|\sigma|=n\right\}$. Then, $a_{1}\leq\dfrac{1}{1+C_{3}}$ and $a_{n+1}\leq\dfrac{a_{n}}{1+C_{3}a_{n}},n\geq 1$. Hence,
$a_{n}\leq\frac{1}{nC_{3}+1}\leq\frac{1}{nC_{3}}.$
Proposition 2.12 (Weak version of regularity).
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are both LF systems and let $\varphi$ be the solution of (1.1). Then, there exists a constant $C>0$ such that for every $y\in(0,1)$ and $r>0$ satisfying that $0\leq y-r<y+r\leq 1$,
$\mu_{\varphi}\left([y-r,y+r]\right)\geq\exp(-C/r).$
This might be regarded as a weak version of the regularity of $\mu_{\varphi}$ in the sense of Ullman-Stahl-Totik. See [22, Criterion 1.1]. However, the following proof does not yield the regularity.
Proof.
We apply Lemma 2.10 to $\{g_{i}\}_{i}$ and Lemma 2.11 to $\{f_{i}\}_{i}$. Let $C$ be the constant appearing in Lemma 2.11. Take $y\in(0,1)$ and $r>0$ such that $0\leq y-r<y+r\leq 1$ arbitralily. Let $n$ be a natural number such that $\frac{C}{n}<r\leq\frac{C}{n-1}$. Then, by the compatibility condition and Lemma 2.11, there exist $i_{1},\cdots,i_{n}\in\mathcal{I}_{N}$ such that
$[f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)]\subset[y-r,y+r].$
Hence, by (2.3),
$\mu_{\varphi}([y-r,y+r])\geq g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0).$
Let $c$ be the constant appearing in Lemma 2.10. By Lemma 2.10,
$\mu_{\varphi}([y-r,y+r])\geq c^{n}\geq c^{1+C/r}.$
3. Singularity
In this section, we consider (1.1) under the setting that $\{g_{i}\}_{i}$ is an LF-system on $[0,1]$, and, $\{f_{i}\}_{i}$ is a weakly contractive D-system on $[0,1]$. By Lemma 2.8, both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are D-systems. Hence, by Lemma 2.3, there exists a unique continuous strictly increasing solution $\varphi$ of (1.1).
We consider singularity for the solution $\varphi$. We will give sufficient conditions for $\dim_{H}\mu_{\varphi}<1$. It is known that if $\dim_{H}\mu_{\varphi}<1$, then, $\varphi$ is a singular function on $[0,1]$, specifically, $\varphi$ is a strictly increasing and continuous function on $[0,1]$ such that its derivative is equal to zero almost everywhere.
The main purpose of this section is to extend [27, Theorem 43] to the case that $f_{i}$ can be non-linear. In [27], the author dealt with many objects other than the conjugate functional equation, so the arguments were not specialized for the functional equation. Some parts dealing with the functional equation in [27] are somewhat sketchy, so we give more details here. We introduce some new notations and the formulations of our main results and some of the arguments in the proofs are different from [27, Theorem 43]. These changes will make arguments more transparent.
We assume that an LF-system $\{g_{i}\}_{i}$ is defined by $g_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers satisfying the conditions (1), (2) and (3) in Definition 2.7. In this section, we assume $d_{i}=1$ for each $i$.
Let
$\alpha:=\min\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\middle|1\leq i\leq N-1\right\},$
where we delete [^2] This is different from [27]. $\frac{c_{0}}{1-a_{0}}$ if $a_{0}=1$. Let
$\beta:=\max\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\middle|1\leq i\leq N-1\right\},$
where $\dfrac{c_{0}}{1-a_{0}}:=+\infty$ if $a_{0}=1$.
Let $Y:=[\alpha,\beta]$. Let
$G_{i}(y):=\frac{(a_{i}-b_{i}c_{i})(y+1)}{(b_{i}y+1)((a_{i}+b_{i})y+(c_{i}+1))}=\frac{(b_{i+1}-b_{i})(y+1)}{(b_{i+1}y+1)(b_{i}y+1)},\ y\in[\alpha,\beta].$
Let
$H_{i}(y):=\frac{a_{i}y+c_{i}}{b_{i}y+1},\ y\in[\alpha,\beta].$
Let $G_{i}(+\infty)$ and $H_{i}(+\infty)$ be the limits. Specifically,
$G_{0}(+\infty):=1,G_{i}(+\infty):=0,H_{0}(+\infty):=+\infty,H_{i}(+\infty):=\frac{a_{i}}{b_{i}},1\leq i\leq N-1.$
The following is the same as [27, Lemma 27], so we will not repeat the proof.
Lemma 3.1.
(i) $0\in Y\subset[-1,+\infty]$.
(ii) Each $H_{i}$ is well-defined on $Y$.
(iii) $\alpha\leq H_{i}(\alpha)\leq H_{i}(\beta)\leq\beta$, that is, $H_{i}(Y)\subset Y$.
(iv) $\sum_{i\in\Sigma_{N}}G_{i}(y)=1,\ y\in Y.$
By (iii), we can compose $H_{i}$ freely on $Y$. We consider the compactification when $\beta=+\infty$. Then, $Y=[\alpha,\beta]$ is compact, and $G_{i}:Y\to[0,1]$ and $H_{i}:Y\to Y$ are continuous. We remark that $Y,G_{i},H_{i}$ are defined by $(g_{i})_{i}$, not by $(f_{i})_{i}$.
$\textup{Fix}(H_{i})$ denotes the set of fixed points of $H_{i}$ on $[\alpha,\beta]$. This is uniquely determined and
$\textup{Fix}(H_{0})=\left\{\dfrac{c_{0}}{1-a_{0}}\right\},\ \textup{Fix}(H_{i})=\left\{\dfrac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\right\},\ i\geq 1.$
Let
$\mathcal{P}_{N}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}>0,\sum_{i}p_{i}=1\right\}$
and
$\overline{\mathcal{P}_{N}}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}\geq 0,\sum_{i}p_{i}=1\right\}.$
Let
$\mathfrak{p}_{i}:=\left(G_{j}(\textup{Fix}(H_{i}))\right)_{j}\in\overline{\mathcal{P}_{N}},\ i\in\mathcal{I}_{N}.$
Let
$\dim_{H}\mu_{\varphi}:=\inf\left\{\dim_{H}A:A\in\mathcal{B}([0,1]),\mu_{\varphi}(A)=1\right\}.$
The following results extend [27, Theorem 43 (i)] to the case that $\{f_{i}\}_{i}$ may be non-affine.
Theorem 3.2.
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$, where we let $r_{i}:=\textup{Lip}(f_{i})$.
The probability vectors $\mathfrak{p}_{i},i\in\mathcal{I}_{N}$, are determined only by $(\alpha,\beta,\{G_{i}\}_{i},\{H_{i}\}i)$ and date back to $\{g_{i}\}_{i}$. $\{f_{i}\}_{i}$ is not related to this. This contains the affine case that $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i$. The reason why we have introduced $Y,G_{i}$ and $H_{i}$ will be clarified in the proof.
We also deal with the case that |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$. Let the total distance between probability vectors be
$\|{\bf p}-{\bf q}\|:=\sum_{i\in\mathcal{I}_{N}}|p_{i}-q_{i}|,\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
Theorem 3.3.
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$, and $\|{\bf r}-\mathfrak{p}_{0}\|\geq\epsilon$, where we let $r_{i}:=\textup{Lip}(f_{i})$ and ${\bf r}:=(r_{0},\cdots,r_{N-1})$.
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$. If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i$, and $\mathfrak{p}_{0}={\bf q}$, then, the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$. See [27, Theorem 43 (ii)].
3.1. Proofs
We first give notations for symbolic dynamics. Let :=$\mathbb{N}:=\{1,2,\cdots\}$. We define a map $\pi:(\mathcal{I}_{N})^{\mathbb{N}}\to[0,1]$ by
$\pi(i_{1}i_{2}\cdots):=\lim_{k\to\infty}f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0).$
This is well-defined because each $f_{i}$ is a contraction and $\left(f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0)\right)_{k}$ is a Cauchy sequence.
Let $\xi_{j}:(\mathcal{I}_{N})^{\mathbb{N}}\to\mathcal{I}_{N}$ be the projection to the $j$-th coordinate. Let
[…]
Then, by Lemma 2.1,
$\pi\left(I(i_{1},\cdots,i_{k})\right)=\left[f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0),f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(1)\right].$
Hence, by (2.3),
$\mu_{\varphi}\left(\pi(I(i_{1},\cdots,i_{k}))\right)$ $=\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(1)\right)-\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(0)\right)$
$=g_{i_{1}}\circ\cdots\circ g_{i_{k}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{k}}(0).$
For $y\in Y$, let $\nu_{y}$ be a probability measure on $(\mathcal{I}_{N})^{\mathbb{N}}$ such that $\nu_{y}(I(i))=G_{i}(y)$, $i\in\mathcal{I}_{N}$, and,
$\frac{\nu_{y}(I(i_{1},\cdots,i_{n-1},i_{n}))}{\nu_{y}(I(i_{1},\cdots,i_{n-1}))}=G_{i_{n}}\circ H_{i_{n-1}}\circ\cdots\circ H_{i_{1}}(y),\ n\geq 2.$ (3.1)
Since $\sum_{i\in\mathcal{I}_{N}}G_{i}(y)=1$ for each $y\in Y$, the consistency condition holds. Therefore, a probability measure on $(\mathcal{I}_{N})^{\mathbb{N}}$ satisfying (3.1) uniquely exists by the Kolmogorov extension theorem.
For $i\in\mathcal{I}_{N}$, let $F_{i}:(\mathcal{I}_{N})^{\mathbb{N}}\to(\mathcal{I}_{N})^{\mathbb{N}}$ be a map defined by $F_{i}(z)=iz$. Now we see that
$\nu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\nu_{H_{i}(y)}\circ F_{i}^{-1}.$ (3.2)
Let $\mu_{y}:=\nu_{y}\circ\pi^{-1}$. Then, $\mu_{y}$ is a Borel probability measure on $[0,1]$. By $\pi\circ F_{i}=f_{i}\circ\pi$ and (3.2),
$\mu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\mu_{H_{i}(y)}\circ f_{i}^{-1}.$ (3.3)
We recall that $0\in Y=[\alpha,\beta]$. Then, by (3.3),
$\mu_{0}=\sum_{i\in\mathcal{I}_{N}}G_{i}(0)\mu_{H_{i}(0)}\circ f_{i}^{-1}.$
Since $H_{i}(0)\neq 0$ can occur, this does not define an iterated function system. (3.3) is in the framework of a place-dependent iterated function system [9].
By [27, Lemma 30],
$\nu_{0}\circ\pi^{-1}=\mu_{0}=\mu_{\varphi}.$ (3.4)
Let $\epsilon\in(0,1/2)$. We assume that $\epsilon\leq\min_{i}r_{i}\leq\max_{i}r_{i}\leq 1-\epsilon$. Since $\{f_{i}\}_{i}$ is a compatible system, $\sum_{i}r_{i}\geq 1$ and $r_{i}>0$ for each $i$. Let $s\geq 1$ is a constant such that $\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$.
As the following assertion shows, the constant $s$ is close to $1$ if $\delta>0$ is sufficiently small.
Lemma 3.4.
Let $\epsilon_{2}>0$. Then, for $\delta>0$ such that
$\frac{\log N}{\log(N/(1+\delta))}<1+\frac{\epsilon_{2}}{\log 1/\epsilon},$ (3.5)
it holds that $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$ for every choice of $(r_{i})_{i}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$.
Proof.
Let $u_{i}:=r_{i}/\sum_{i}r_{i}$. Then, $\sum_{i}u_{i}=1$. We see that $\sum_{i}r_{i}^{s}=1$ if and only if $\sum_{i}u_{i}^{s}=(\sum_{i}r_{i})^{-s}$. Hence, if $\sum_{i}r_{i}\leq 1+\delta$, then,
$\frac{\log(\sum_{i}u_{i}^{s})}{s}\leq-\log(1+\delta).$
$\sum_{i}u_{i}^{s}$ takes its minimum at $u_{0}=\cdots=u_{N-1}=\dfrac{1}{N}$. We see this by the method of Lagrange multipliers. The minimum is $N^{1-s}\leq 1$. Hence, $-\dfrac{s-1}{s}\log N\geq-\log(\sum_{i}r_{i})$, which is equivalent to $s<\dfrac{\log N}{\log(N/(1+\delta))}$. Therefore, if (3.5) holds, then, $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$. ∎
Let
$R_{n}({\bf q};z):=\frac{\nu_{0}(I_{n}(z))}{q_{\xi_{1}(z)}\cdots q_{\xi_{n}(z)}},\ \ {\bf q}=(q_{i})_{i}\in\mathcal{P}_{N},z\in(\mathcal{I}_{N})^{\mathbb{N}}.$
By a measure-theoretic argument, we can show that
Lemma 3.5.
Let $\widetilde{q}_{i}:=r_{i}^{s}>0$ and ${\bf\widetilde{q}}:=(\widetilde{q}_{i})_{i}\in\mathcal{P}_{N}$. If there exists $\epsilon_{2}>0$ such that
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf\widetilde{q}};z)}{n}\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$ (3.6)
then,
$\dim_{H}\mu_{\varphi}\leq s-\frac{\epsilon_{2}}{\log 1/\epsilon}.$
We omit the proof of this assertion. See the proof of [27, Lemma 7]. (3.4) plays an important role.
By Lemmas 3.4 and 3.5, we obtain that
Proposition 3.6.
Assume that there exists $\epsilon_{2}>0$ such that for every ${\bf q}\in\mathcal{P}_{N}$,
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$ (3.7)
Then, for $\delta=\delta(N,\epsilon,\epsilon_{2})>0$ satisfying (3.5), $\dim_{H}\mu_{\varphi}<1$.
Now it suffices to establish (3.7). We give some definitions for the relative entropy. Let
$s_{N}\left({\bf p}|{\bf q}\right):=\sum_{i\in\mathcal{I}_{N}}p_{i}\log\frac{q_{i}}{p_{i}},\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
If $p_{i}=0$, then, $p_{i}\log\frac{q_{i}}{p_{i}}:=0$. $s_{N}$ is continuous on $\overline{\mathcal{P}_{N}}$. Then, by Jensen’s inequality, $s_{N}({\bf p}|{\bf q})\leq 0$. It holds that $s_{N}\left({\bf p}|{\bf q}\right)=0$ if and only if ${\bf p}={\bf q}$.
For $y\in Y,z\in(\mathcal{I}_{N})^{\mathbb{N}}$, let
$p_{j}(y;z):=\left(G_{i}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(y)\right)_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}},\ j\geq 1.$
and
$p_{0}(y;z):=(G_{i}(y))_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}}.$
By a martingale argument, we can show that
Proposition 3.7.
It holds that for every ${\bf q}\in\mathcal{P}_{N}$,
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}=\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}(p_{j}(0;z)|{\bf q}),\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
We omit the proof of this assertion. See the proof of [27, Proposition 6]. It is important to assume that ${\bf q}\in\mathcal{P}_{N}$, not ${\bf q}\in\overline{\mathcal{P}_{N}}$.
For $\epsilon\in(0,1)$, let $\overline{\mathcal{P}_{N}(\epsilon)}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}\geq\epsilon,\sum_{i}p_{i}=1\right\}$. Then, $\overline{\mathcal{P}_{N}(\epsilon)}\subset\mathcal{P}_{N}\subset\overline{\mathcal{P}_{N}}$.
Recall Lemma 3.4. If
$\frac{\log N}{\log(N/(1+\delta))}\leq 1+\frac{\log 2}{\log(1/\epsilon)},$
then, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. In other words, ${\bf\widetilde{q}}$ is apart from the boundary $\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$
Assume that
Proposition 3.8.
If |$|\{\mathfrak{p}_{i}\}_{i}|\geq 2$, then, there exists $\epsilon_{2}>0$ such that for every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
Theorem 3.2 follows from Propositions 3.6, 3.7 and 3.8. We now show Proposition 3.8. We give a strategy of the proof. By the definition, $s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq 0$. Hence, it suffices to show the number of $j$ satisfying that $s_{N}\left(p_{j}(0;z)|{\bf q}\right)$ is strictly negative has positive density in $\mathbb{N}$.
We first give some further definitions for quantities of the relative entropy.
Lemma 3.9.
Let
$V\left({\bf q};\epsilon\right):=\max\left\{s_{N}\left({\bf p}|{\bf q}\right)\middle|{\bf p}\in\overline{\mathcal{P}_{N}},\|{\bf p}-{\bf q}\|\geq\epsilon\right\},\ \epsilon>0.$
Then,
(i) $V\left({\bf q};\epsilon\right)<0$.
(ii) $V\left({\bf q};\epsilon\right)$ is decreasing with respect to $\epsilon$.
(iii) $V\left({\bf q};\epsilon\right)$ is right-continuous with respect to $\epsilon$.
(iv) $V\left({\bf q};\epsilon\right)$ is continuous with respect to ${\bf q}$.
Proof.
(i) and (ii) are easy to see. (iii) Assume that for some $\epsilon_{0}\in(0,1/3)$, $V\left({\bf q};\epsilon\right)$ is not right-continuous. Then, $V\left({\bf q};\epsilon_{0}\right)>\lim_{\epsilon\to\epsilon_{0}+0}V\left({\bf q};\epsilon\right)$. Assume that $V\left({\bf q};\epsilon_{0}\right)=s_{N}\left({\bf p}|{\bf q}\right)$. If $\|{\bf p}-{\bf q}\|>\epsilon_{0}$, then, this cannot occur. Assume $\|{\bf p}-{\bf q}\|=\epsilon_{0}$. Then, there exists a sequence $\{{\bf p}^{(n)}\}_{n}$ in $\overline{\mathcal{P}_{N}}$ such that $\|{\bf p}^{(n)}-{\bf q}\|\geq\epsilon_{0}+N^{-n}$ for each $n$ and $\lim_{n\to\infty}\|{\bf p}^{(n)}-{\bf q}\|=\epsilon_{0}$. Hence,
$\liminf_{n\to\infty}V\left({\bf q};\epsilon_{0}+N^{-n}\right)\geq\lim_{n\to\infty}s_{N}\left({\bf p}^{(n)}|{\bf q}\right)=s_{N}\left({\bf p}|{\bf q}\right)=V\left({\bf q};\epsilon_{0}\right).$
This is a contradiction.
(iv) follows from (iii) and
$V\left({\bf q};\epsilon+\|{\bf p}-{\bf q}\|\right)\leq V\left({\bf p};\epsilon\right)$
By Lemma 3.9, we can let
$V(\epsilon):=\max_{{\bf q}\in\overline{\mathcal{P}_{N}}}V\left({\bf q};\epsilon\right)<0.$
We now give a quantity defined by $Y,\{G_{i}\}_{i}$ and $\{H_{i}\}_{i}$.
Lemma 3.10.
Let
$W_{i}\left({\bf q},y\right):=\max\left\{\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|,\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|\right\},\ {\bf q}\in\overline{\mathcal{P}_{N}},y\in Y,i\in\mathcal{I}_{N}.$
Then,
(i) $W_{i}({\bf q},y)$ is continuous on $\overline{\mathcal{P}_{N}}\times Y$.
(ii) $W_{i}({\bf q},y)=0$ if and only if ${\bf q}=\mathfrak{p}_{i}$ and $y=\textup{Fix}(H_{i})$.
This assertion is easily shown. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then, $s_{N}\left((G_{k}(y))_{k}|{\bf q}\right)<0$. However, in our setting, we cannot exclude the possibility that $|G_{k}(y)-q_{k}|=0$ for every $k\in\mathcal{I}_{N}$. Assume that |$\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|$ is close to 00, then, $\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|$ is strictly positive. Hence, $W_{i}\left({\bf q},y\right)>0$.
The following corresponds to [27, Lemma 8]. Informally speaking, if $W_{i}({\bf q},y)$ is positive uniformly to $y$, then, a sum of two successive terms of $s_{N}$ is negative.
Lemma 3.11.
Let ${\bf q}\in\mathcal{P}_{N}$. Let $i_{1}\in\mathcal{I}_{N}$. Assume that for some $\epsilon_{1}>0$,
$\inf_{y\in Y}W_{i_{1}}\left({\bf q},y\right)\geq\epsilon_{1}.$ (3.8)
Assume that $X_{j+1}(z)=i_{1}$. Then,
$s_{N}\left(p_{j}(0;z)|{\bf q}\right)+s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)\leq V(\epsilon_{1})<0.$
Proof.
The proof is essentially the same as the proof of [27, Lemma 8] [^3] It contains a typo.. Since $X_{j+1}(z)=i_{1}$,
$p_{j}(0;z)=\left(G_{k}(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))\right)_{k\in\mathcal{I}_{N}},$
and
$p_{j+1}(0;z)=\left(G_{k}(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))\right)_{k\in\mathcal{I}_{N}}.$
Let ${\bf q}=(q_{k})_{k}$. Then, by (3.8), either
$\left|G_{k_{1}}(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))-q_{k_{1}}\right|\geq\epsilon_{1},\textup{ for some }k_{1}\in\mathcal{I}_{N},$
or,
$\left|G_{k_{2}}(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))-q_{k_{2}}\right|\geq\epsilon_{1},\textup{ for some }k_{2}\in\mathcal{I}_{N},$
holds. If the former holds, then, $s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$. If the latter holds, then, $s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$. Now it suffices to recall that $s_{N}$ takes non-positive values. ∎
Let :=|$\ell:=|\{\mathfrak{p}_{i}\}_{i}|\geq 2$. We assume that $\{\mathfrak{p}_{i}\}_{i}=\{\mathfrak{p}_{j_{k}}:1\leq k\leq\ell\}$ where $10\leq j_{1}<\cdots<j_{\ell}\leq N-$ 1. Let $\delta_{0}:=\min\{\|\mathfrak{p}_{j_{k_{1}}}-\mathfrak{p}_{j_{k_{2}}}\|:k_{1}\neq k_{2}\}>0$. Let $B_{j}:=\left\{{\bf p}\in\overline{\mathcal{P}_{N}}:\|{\bf p}-\mathfrak{p}_{j}\|\leq\delta_{0}/4\right\}$. $B_{j}=B_{k}$ can occur even if $j\neq k$. Define a coloring map $\mathscr{A}:\overline{\mathcal{P}_{N}}\to\mathcal{I}_{N}$ by
$\mathscr{A}\left({\bf p}\right):=\begin{cases}j_{2}&{\bf p}\in B_{j_{1}}\\ \vdots&\\ j_{\ell}&{\bf p}\in B_{j_{\ell-1}}\\ j_{1}&{\bf p}\in B_{j_{\ell}}\\ 0&\textup{ otherwise}\end{cases}.$
Under these definitions, we see that
Lemma 3.12.
$\widetilde{\epsilon_{1}}:=\inf\left\{W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\middle|{\bf q}\in\overline{\mathcal{P}_{N}},y\in Y\right\}>0.$
Proof.
Fix ${\bf q}\in\overline{\mathcal{P}_{N}}$. By Lemma 3.10,
$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{2}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|.$
If $(G_{j}(y))_{j}\notin B_{\mathscr{A}({\bf q})}$, then, $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}=\left(G_{j}(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)}))\right)_{j}$ and hence, $y$ is not contained in an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$. Since $G_{j}$ and $H_{\mathscr{A}\left({\bf q}\right)}$ are continuous, $\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$ is continuous with respect to $y$ on $Y$. Hence,
$\inf_{(G_{j}(y))_{j}\notin B_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
$\geq\inf_{y\notin J_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|=:\epsilon_{\mathscr{A}\left({\bf q}\right)}>0.$
Assume that $(G_{j}(y))_{j}\in B_{\mathscr{A}({\bf q})}$. Since ${\bf q}\notin B_{\mathscr{A}({\bf q})}$,
$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{N}\sum_{j}|G_{j}(y)-q_{j}|\geq\frac{\delta_{0}}{4N}.$
Hence,
$\inf_{y\in Y}W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\min\left\{\frac{\delta_{0}}{4N},\epsilon_{j_{k}}\middle|1\leq k\leq\ell\right\}=:\tilde{\epsilon_{1}}.$
We discuss the frequency of occurrence of some elements in $\mathcal{I}_{N}$ in $(\xi_{n})_{n}$.
Lemma 3.13.
Let $I\subset\mathcal{I}_{N}$ be a non-empty set. Let $c_{I}:=\inf_{y\in Y}\sum_{i\in I}G_{i}(y)$. Then,
$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n:\xi_{j}(z)\in I\}|}{n}\geq c_{I},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
This corresponds to [27, Lemma 9], but the formulation differs.
Proof.
If $c_{I}=0$, then, the assertion is obvious. Assume $c_{I}>0$. We denote the expectation with respect to $\nu_{0}$ by $E^{\nu_{0}}$. Let $\mathcal{F}_{m}:=\sigma(\xi_{1},\cdots,\xi_{m})$ for $m\geq 1$ and $\mathcal{F}_{0}$ be the trivial $\sigma$-algebra on $(\mathcal{I}_{N})^{\mathbb{N}}$. We denote the indicator function of $A\subset(\mathcal{I}_{N})^{\mathbb{N}}$ by ${\bf 1}_{A}$.
Let $M_{0}:=0$ and
$M_{n}-M_{n-1}:={\bf 1}_{\{\xi_{n}\in I\}}-E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{n}\in I\}}|\mathcal{F}_{n-1}\right],n\geq 1.$
Then, $\{M_{n},\mathcal{F}_{n}\}_{n}$ is a martingale.
By (3.1), $E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{n}\in I\}}|\mathcal{F}_{n-1}\right]\geq c_{I}$, $\nu_{0}$-a.s. Therefore, for every $\eta\in(0,1)$,
$\nu_{0}\left(\frac{|\{1\leq j\leq n:\xi_{j}\in I\}|}{n}\leq(1-\eta)c_{I}\right)\leq\nu_{0}\left(M_{n}\leq-n\eta c_{I}\right).$
We remark that $|M_{n}-M_{n-1}|\leq 1$ for each $n\geq 1$. By Azuma’s inequality ([1], [31, E14.2]),
$\nu_{0}\left(M_{n}\leq-n\eta c_{I}\right)\leq\exp\left(-\frac{n\eta^{2}c_{I}^{2}}{2}\right).$
Using $c_{I}>0$ and Borel-Cantelli’s lemma,
$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n:\xi_{j}(z)\in I\}|}{n}\geq(1-\eta)c_{I},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
The assertion follows if $\eta\to+0$. ∎
We proceed to the proof of Proposition 3.8. The proof is divided into two cases. The following is easily shown.
Lemma 3.14.
(i) $a_{0}\leq 1$ and $b_{N-1}+c_{N-1}\geq 0$.
(ii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then, $\alpha=-1,\beta=+\infty$.
(iii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$, then, $\alpha>-1,\beta=+\infty$.
(iv) If $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$, then, $\alpha=-1,\beta<+\infty$.
(v) If $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$, then, $\alpha>-1,\beta<+\infty$.
(vi) If $a_{0}=1$, then, $\textup{Fix}(H_{0})=\{+\infty\}$ and $\mathfrak{p}_{0}=(1,0,\cdots,0)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
(vii) If $b_{N-1}+c_{N-1}=0$, then, $\textup{Fix}(H_{N-1})=\{-1\}$ and $\mathfrak{p}_{N-1}=(0,\cdots,0,1)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
$\alpha=-1$ or $\beta=+\infty$ occurs according to the values of $a_{0}$ and $b_{N-1}+c_{N-1}$.
Case 1. We first deal with the case that $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$.
In this case, we do not need to assume ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
Proof of Proposition 3.8, Case 1.
By the assumption that $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$ and Lemma 3.14 (v), $-1<\alpha\leq\beta<+\infty$. Hence,
$10<\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)\leq\sup_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)<$ 1 .
Fix ${\bf q}\in\overline{\mathcal{P}_{N}}$. Let $\widetilde{c}:=\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)$. Then, $c_{\{\mathscr{A}({\bf q})\}}\geq\widetilde{c}>0$ and by Lemma 3.13 for $I=\{\mathscr{A}({\bf q})\}$,
$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}{\bf 1}_{\{\xi_{k}=\mathscr{A}({\bf q})\}}(z)\geq\widetilde{c},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
The number $i_{1}\in\mathcal{I}_{N}$ in Lemma 3.11 will be chosen according to the position of ${\bf q}$. By applying Lemma 3.11 to $i_{1}=\mathscr{A}({\bf q})$ and $\epsilon_{1}=\widetilde{\epsilon_{1}}$ where $\widetilde{\epsilon_{1}}$ is the constant in Lemma 3.12, we see that if $X_{i+1}(z)=\mathscr{A}({\bf q})$, then,
$s_{N}\left(p_{i}(0;z)|{\bf q}\right)+s_{N}\left(p_{i+1}(0;z)|{\bf q}\right)\leq V(\widetilde{\epsilon_{1}})<0.$
Now the assertion follows if we let $\epsilon_{2}:=\dfrac{\widetilde{c}\ V(\widetilde{\epsilon_{1}})}{2}$. ∎
Case 2. We now deal with the case that $a_{0}=1$ or $b_{N-1}+c_{N-1}=0$.
If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then, $c_{\{0\}}=c_{\{N-1\}}=0$, so we consider $c_{\{0,N-1\}}$. In the following, we do not use Lemma 3.12 or the assumption that |$|\{\mathfrak{p}_{i}\}_{i}|\geq 2$. On the other hand, we assume ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
Proof of Proposition 3.8, Case 2.
Case 2.1. We first consider the case that $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$.
By Lemma 3.14, $\alpha>-1$ and $\beta=+\infty$, $H_{0}(y)=y+c_{0}$, and $\textup{Fix}(H_{0})=\{+\infty\}$. $\mathfrak{p}_{0}=(1,0,\cdots,0)\notin\mathcal{P}_{N}$. On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\geq 1$.
Since $G_{0}(+\infty)=G_{0}(H_{0}(+\infty))=1$ and $q_{0}<1-\epsilon/2$,
$\inf_{y\in Y}W_{0}({\bf q},y)>c_{0}(\epsilon/2)>0,$
where the lower bound $c_{0}(\epsilon/2)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
Since $\alpha>-1$, $\inf_{y\in Y}G_{0}(y)\geq c>0$. By using Lemma 3.11 and Lemma 3.13 for $I=\{0\}$, we have the assertion.
Case 2.2. We secondly consider the case that $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$.
We recall that $\alpha=-1$ and $\beta<+\infty$, By Lemma 3.14, $H_{N-1}(y)=\dfrac{(1-2b_{N-1})y+b_{N-1}}{1-b_{N-1}}$, and $\textup{Fix}(H_{N-1})=\{-1\}$. $\mathfrak{p}_{N-1}=(0,\cdots,0,1)\notin\mathcal{P}_{N}$. On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\leq N-2$.
Since $G_{N-1}(-1)=G_{N-1}(H_{N-1}(-1))=1$ and $q_{N-1}<1-\epsilon/2$,
$\inf_{y\in Y}W_{N-1}({\bf q},y)>c_{N-1}(\epsilon/2)>0,$
where the lower bound $c_{N-1}(\epsilon)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. Since $\beta<+\infty$, $\inf_{y\in Y}G_{N-1}(y)\geq c>0$. The rest of the argument is the same as in the above case.
Case 2.3. We finally consider the case that $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$.
Since $\textup{Fix}(H_{0})=\{+\infty\}$, $\textup{Fix}(H_{N-1})=\{-1\}$, $q_{0}<1-\epsilon/2$ and $q_{N-1}<1-\epsilon/2$, it holds that
$\inf_{y\in Y}W_{0}({\bf q},y)>c_{0}(\epsilon/2)>0$
and
$\inf_{y\in Y}W_{N-1}({\bf q},y)>c_{N-1}(\epsilon/2)>0.$
Since
$G_{0}(y)+G_{N-1}(y)=\frac{b_{1}b_{N-1}y^{2}+2b_{1}y1+b_{1}-b_{N-1}}{(b_{1}y+1)(b_{N-1}y+1)},$
it holds that
$\inf_{y\in Y}G_{0}(y)+G_{N-1}(y)>0.$
By using Lemma 3.11 and Lemma 3.13 for $I=\{0,N-1\}$, we have the assertion. ∎
Now the proof of Theorem 3.2 completes.
Remark 3.15.
(i) In the proof of [27, Theorem 43], the details of Case 2 were not given.
(ii) Assume that $\{g_{i}\}_{i}$ are affine functions such that $g_{i}(x)=b_{i}x+(b_{0}+\cdots+b_{i-1})$. Since $Y=\{0\}$, there is no need to introduce $H_{i}$, and $W_{i}({\bf q},y)=\max_{k}|G_{k}(y)-q_{k}|=\max_{k}|b_{k}-q_{k}|$. However, $\mathscr{p}_{0}=\cdots=\mathscr{p}_{N-1}=(b_{i})_{i}$, and hence this is not in the scope of Theorem 3.2. This case is considered in Theorem 3.3.
We now proceed to the proof of Theorem 3.3.
By (3.6), if we take $\delta>0$ such that
$\frac{\log N}{\log(N/(1+\delta))}\leq 1+\frac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)},$ (3.9)
then, $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$.
Assume that $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$. Since $\min_{i}r_{i}\geq\epsilon$, $\sum_{i}|r_{i}-r_{i}^{s}|\leq\epsilon^{1-s}-1\leq\epsilon/2$. Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$, where we let $\widetilde{q}_{i}:=r_{i}^{s}$ for each $i$ and ${\bf\widetilde{q}}:=(\widetilde{q}_{i})_{i}$. Since $\epsilon\leq 1/2$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\epsilon/2$ for each $i$. Hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
By Lemma 3.10 (i) and (ii), it holds that
$\epsilon_{3}:=\inf\left\{W_{i}({\bf q},y)\middle|i\in\mathcal{I}_{N},y\in Y,{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2\right\}>0.$
By Lemma 3.11, we see that for every $j\geq 0$, every $z\in(\mathcal{I}_{N})^{\mathbb{N}}$, and every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$ with $\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$,
$s_{N}\left(p_{j}(0;z)|{\bf q}\right)+s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{3}\right).$
By Lemma 3.9, we can let
$V_{\epsilon}(\epsilon_{3}):=\max\left\{V\left({\bf q};\epsilon_{3}\right)\middle|{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2\right\}<0.$
Then,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\frac{V_{\epsilon}(\epsilon_{3})}{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
Assume that $\delta>0$ satisfies not only (3.9) but also (3.5) for $\epsilon_{2}:=V_{\epsilon}(\epsilon_{3})/2$. Then, by Propositions 3.6 and 3.7, $\dim_{H}\mu_{\varphi}<1$.
Remark 3.16.
(i) We do not need to use Lemma 3.13. For every small $\epsilon>0$ and every $i\in\mathcal{I}_{N}$,
$\min\left\{W_{i}({\bf q},y)\middle|y\in Y,{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}\right\}=0,$
by Lemma 3.10 (ii).
(ii) The case that both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are affine is easily handled by the dimension formula. However, if some of the functions $\{f_{i},g_{i}\}_{i}$ are non-affine, then, it is difficult to judge the singularity by a generalized dimension formula in [9].
4. Example
We give an example that $\{g_{i}\}_{i}$ is an LF system and $\{f_{i}\}_{i}$ is non-affine and the solution $\varphi$ is singular. The result depends on the quantitative estimates in Section 3.
Let $N=2$. Then,
$G_{0}(y)=\frac{b_{1}(y+1)}{b_{1}y+1},\ H_{0}(y)=a_{0}y+c_{0},\ H_{1}(y)=\frac{(1-b_{1}+c_{1})y+c_{1}}{b_{1}y+1},$
Let
$g_{0}(x):=\frac{x}{-x+6},\ g_{1}(x):=\frac{3x+1}{5-x}.$
Then, $g_{0}$ is contractive, and $g_{1}$ is weakly contractive but not contractive. It holds that $a_{0}=1/6$, $c_{0}=-1/6$, $a_{1}=3/5$, $b_{1}=1/5$, and $c_{1}=-1/5$. Hence,
$\alpha=\min\{0,-1/5,-1\}=-1,\beta=\max\{0,-1/5,-1\}=0.$
Hence,
$\sup_{y\in[-1,0]}G_{0}(y)=1-\inf_{y\in[-1,0]}G_{1}(y)=\frac{1}{5}.$
Let $\mathcal{P}[4/5]:=\left\{(p_{0},p_{1})\in\mathcal{P}_{2}:p_{0}\geq 4/5\right\}$. Let ${\bf q}=(q_{0},q_{1})\in\mathcal{P}[4/5]$. Then,
$|G_{0}(y)-q_{0}|=|G_{1}(y)-q_{1}|\geq\frac{3}{5},\ y\in[-1,0].$
Hence,
$s_{2}\left(p_{j}(0,z)|{\bf q}\right)\leq V\left({\bf q};\frac{6}{5}\right),\ j\geq 0,z\in(\mathcal{I}_{N})^{\mathbb{N}}.$ (4.1)
Since $s_{2}({\bf p}|{\bf q})=s_{2}\left((p_{0},1-p_{0})|(q_{0},1-q_{0})\right)$ is increasing with respect to $p_{0}$ on $[0,q_{0}]$,
$V\left({\bf q};\frac{6}{5}\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|(q_{0},1-q_{0})\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|\left(\frac{4}{5},\frac{1}{5}\right)\right)\leq-\frac{6}{5}\log 2\leq-0.83.$
By Proposition 3.6, it holds that
$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.83,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\mathcal{P}[4/5]$.}$ (4.2)
Let $\epsilon\in(0,1/2)$. Let $\{f_{0},f_{1}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1$. Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}=1$. Assume that $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$. Then, by (4.2) and Lemma 3.5,
$\dim_{H}\mu_{\varphi}\leq s-\frac{0.83}{\log(1/\epsilon)}.$
We see that $\dfrac{0.83}{\log 10}\geq 0.36$ and $\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$. Hence, (3.5) holds for $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$. By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, $r_{0}+r_{1}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$, then,
$\dim_{H}\mu_{\varphi}\leq s-\frac{0.83}{\log(1/\epsilon)}\leq\dfrac{\log 2}{\log 2-\log 1.2}-\frac{0.83}{\log 10}<1.$
Now we can let $\{f_{0},f_{1}\}$ be a compatible system such that $0.85\leq r_{0}\leq 0.9$ and $0.1\leq r_{1}\leq 0.3$. Then, $r_{0}+r_{1}\leq 1.2$. By the compatibility condition, $r_{0}+r_{1}\geq 1$. The maps $f_{0}$ and $f_{1}$ may not be affine. Since $s\leq\frac{\log 2}{\log 2-\log 1.2}\leq 1.357$, $r_{0}^{s}\geq(0.85)^{1.357}\geq 0.8$. Hence, $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$.
Remark 4.1.
We see that $a_{0}<1$ and $b_{1}+c_{1}=0$. Hence, this case corresponds to Case 2.2. However, we do not need to apply Lemma 3.11 since we have the upper bound for the relative entropy in (4.1).